OpenStudy (anonymous):

Solve for x: -5(x - 3) = 30

OpenStudy (anonymous):

divide both sides by \(-5\) and get \[x-3=-6\] then add 3

OpenStudy (anonymous):

you get \(x=-3\) and we can check by computing \[-5(-3-3)=-5\times (-6)=30\checkmark\]

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

yw